In crystal field theory, in octahedral complexes why does Eg orbital increase in energy by 3/5(delta), why not 2/5(delta)?

Grade:11

1 Answers

Aditi Bhatnagar
12 years ago

Hi Abhinav

There are 2 eg orbitals while 3 t2g orbitals. so each  eg orbital raised by 3/5 (2*3/5=6/5) and each t2g orbital decreased by -3/5 (2*-3/5=-6/5) so that net energy i.e. energy of barry centre remains zero.

Hope that helps!